OpenStudy (anonymous):

find the x and y intercepts for f(x)=x^2-8x+1

OpenStudy (anonymous):

For x-intercepts, set x = 0. You get f(x) = 1. For y-intercepts, set f(x) = y = 0. Use the quadratic formula or factor/complete the square.

OpenStudy (anonymous):

plug x=0 solve fo y\[y=0^{2}-8(0)+1\]\[y=1\] plug y=0 solve for x \[0=x^{2}-8x+1\] \[\frac{-b \pm \sqrt{b ^{2}-(4\times a \times c} }{2\times a}\] a=1 b=-8 c=1 can u slove for x?

OpenStudy (anonymous):

x=4+or-sqrt3?

OpenStudy (anonymous):

\[\frac{-(-8)\pm \sqrt{-8^{2}-(4\times 1\times 1)}}{2 \times 1}\]\[\frac{8\pm \sqrt{64-4}}{2}= \frac{8\pm \sqrt{60}}{2}\]\[\frac{8\pm2\sqrt{15}}{2}=4\pm \sqrt{15}\]

OpenStudy (anonymous):

oops i added it.

OpenStudy (anonymous):

If I deserve medal then give me one

OpenStudy (anonymous):

T Y